let x be a random variable with E(x)=1 and e[x(x-1)]=4. find var(x)



1
Expert's answer
2022-02-10T04:08:30-0500

We are given that,

"E(x)=1" and "E(x(x-1))=4"

We can write "E(x(x-1))" as,

"E(x(x-1))=E(x^2-x)=E(x^2)-E(x)=4", but "E(x)=1". So,"E(x^2)-E(x)=E(x^2)-1=4\implies E(x^2)=5"  

Now,

"var(x)=E(x^2)-(E(x))^2=5-(1)^2=5-1=4"

Therefore, var(x)=4.